For the problem below, what is the initial volume?

Edelman and Gronkowski want to play a game of beach volleyball. If the beach ball has a volume of 400 L at a temperature of 500 K, what will the temperature of the balloon be if the volume decreases to 176 L?


1
Expert's answer
2022-03-24T02:42:02-0400

Using Charles' law

(V1/T1) = (V2/T2)

(400/500) = (176/T2)

T2 = (176*500)/400

T2 = 220K